# GIST: Greedy Independent Set Thresholding for Diverse Data Summarization

### v1

@@ 3 Algorithm
Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of the GIST algorithm.
@@
Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$. There are two cases: $d^{*}<\varepsilon d_{\max}$ or $d^{*}\geq\varepsilon d_{\max}$. For the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation. To see this, observe that
$\textnormal{OPT}\geq(1-\alpha)d_{\max}>(1-\alpha)\cdot\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>(1-\alpha)d^{*}.$
The sum of the $k$ heaviest points upper bounds $w(S^{*})$, so it follows that
$\textnormal{ALG}\geq\alpha\frac{1}{k}w(S^{*})=\textnormal{OPT}-(1-\alpha)d^{*}>(1-\varepsilon)\textnormal{OPT}.$
Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. It follows that GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
@@
$f(T)$ $\geq\alpha\frac{1}{k}w(S^{*})+(1-\alpha)d\geq\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\frac{d^{*}}{2(1+\varepsilon)}.$ (4)
@@
$\textnormal{ALG}\geq(1-\alpha)d_{\max}\geq(1-\alpha)d^{*}.$ (5)
Putting Equation 4 and (5) together, for any $p\in[0,1]$, it holds that
ALG $\geq p\left[\alpha\frac{1}{k}w(S^{*})+(1-\alpha)\frac{d^{*}}{2(1+\varepsilon)}\right]+(1-p)(1-\alpha)d^{*}$
$=p\cdot\alpha\frac{1}{k}w(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot(1-\alpha)d^{*}.$
To maximize the approximation ratio as $\varepsilon\rightarrow 0$, solve $p=1-p/2$ to get $p=2/3$. It follows that
ALG $\geq\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot(1-\alpha)d^{*}$ (6)
$=\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot(1-\alpha)d^{*}$
$\geq\frac{2}{3}\cdot\alpha\frac{1}{k}w(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot(1-\alpha)d^{*}$
$\geq\frac{2}{3}\left(1-\varepsilon\right)\cdot\alpha\frac{1}{k}w(S^{*})+\frac{2}{3}\left(1-\varepsilon\right)\cdot(1-\alpha)d^{*}$
$\geq\left(\frac{2}{3}-\varepsilon\right)\cdot\textnormal{OPT},$
@@
Remark 3.3.
We can modify GIST to achieve an exact $\nicefrac{{2}}{{3}}$-approximation by trying all pairwise distance thresholds in $D=\{\textnormal{dist}(u,v)/2:u,v\in V\}$. This requires $O(n^{2})$ calls to GreedyIndependentSet compared to $O(\log_{1+\varepsilon}(1/\varepsilon))$, but otherwise the same analysis holds for $\varepsilon=0$.
Finally, we show how to implement GIST in nearly linear time for low-dimensional Euclidean metrics using cover trees for exact nearest-neighbor queries (Beygelzimer et al., 2006; Izbicki and Shelton, 2015) and a $(1-\varepsilon)$-approximation algorithm for point-set diameters $d_{\max}$ (Barequet and Har-Peled, 2001; Imanparast et al., 2018).
Theorem 3.4.
For $V\subseteq\mathbb{R}^{d}$ and Euclidean distance, GIST runs in time $O(n^{3}k+n^{2}d)$. Furthermore, for fixed dimension $d$, the running time can be improved to $O(n\log n\cdot\log_{1+\varepsilon}(1/\varepsilon)\cdot\textnormal{poly}(d)+1/\varepsilon^{d-1})$.
We give the proof in Section A.1 and note that the key property is that cover trees allow us to insert a point into a nearest-neighbor index in $O(c^{12}\log n)$ time and find a nearest neighbor to a query point in $O(c^{6}\log n)$ time, where $c=\Theta(d)$ is the doubling dimension of $d$-dimensional Euclidean space.
4 Hardness of approximation
[…]
  1. 1.
    There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm for general metric spaces.
  2. 2.
    There is no polynomial-time $(\nicefrac{{3}}{{4}}+\varepsilon)$-approximation algorithm for general metric spaces, even if $k\geq|V|/4$.
  3. 3.
    APX-completeness for the Euclidean metric, i.e., there is no polynomial-time approximation scheme (PTAS) for this objective.
4.1 General metric spaces
Our first result builds on the hardness of approximating the size of the maximum clique, and it tightly complements the $\nicefrac{{2}}{{3}}$-approximation guarantee that we can achieve with GIST.
Theorem 4.1.
[…]
@@ 4.1 General metric spaces
We reduce this instance of the maximum clique decision problem to MDDS with objective function (1) as follows. Represent each vertex of graph $G$ with a point in our ground set. The distance between a pair of points is $2$ if there is an edge between their corresponding vertices in $G$, and it is $1$ otherwise.
@@
4.2 Large subsets
In many data curation problems, we are interested in instances of MDDS where $k=\Omega(|V|)$. Even in this restricted case, we can show hardness of approximation.
Theorem 4.2.
[…]
The proof of this theorem uses the vertex cover instance constructed in the proof of Håstad (2001, Theorem 8.1) and is given in Section B.1.
4.3 Euclidean metrics
[…]
Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
Similar to our reduction from maximum clique in Section 4.1 for general metric spaces, this construction uses a node embedding function $h_{g}(v)$ to encode graph adjacency in Euclidean space.
Lemma 4.4.
Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and maximum degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
Theorem 4.5.
The MDDS problem is APX-complete for Euclidean metrics.
We sketch the main idea of our reduction and defer the complete proofs of Lemma 4.4 and Theorem 4.5 to Appendix B. Let $\mathcal{I}(G)$ be the set of independent sets of $G$. A set of embedded nodes $S\subseteq V$ (with at least two nodes) of a graph $G$ with maximum degree $\Delta\leq 3$ has the property:
  • If $S\in\mathcal{I}(G)$: $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
  • If $S\not\in\mathcal{I}(G)$: $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-1/(2(\Delta+1))\leq 1-\nicefrac{{1}}{{8}}$.
This gap between the two cases is at least $\nicefrac{{1}}{{8}}$ (i.e., a universal constant) for all such graphs, and allows us to show that MDDS inherits the APX-complete hardness of bounded-degree max independent set.
5 Experiments
5.1 Warm-up: Synthetic dataset
Setup.
We start with a simple dataset of weighted random points to see the shortcomings of existing methods for MDDS. For a fixed dimension $d$, we generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ and each point is assigned an i.i.d. uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$.
We consider three baseline algorithms: random, simple, and greedy. The random algorithm chooses a random set of $k$ points, permutes them, and returns the best prefix. simple is the $\nicefrac{{1}}{{2}}$-approximation algorithm in Section 2. greedy builds a set $S\subseteq V$ one element at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$. Then it returns the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$, as this sequence of objective values is not necessarily monotone.
Figure 1: Synthetic data with $n=1000$ and $d=64$. Plots of $f(S_{\text{ALG}})$ for $k\in[n]$, $\alpha\in\{0.90,0.95\}$.
Results.
We plot the values of $f(S)$ for our baseline algorithms and GIST for all $k\in[n]$ in Figure 1. We set $\alpha\in\{0.90,0.95\}$ to balance the two competing objective terms (which are easy to optimize in isolation). [^2] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$, so $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$. There is a clear separation in solution quality for each $k\in[n]$ when $\alpha=0.95$. Note that these objective values are not directly comparable as a function of $k$ due to the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective decreasing on average in $k$ for a fixed $\alpha$: mean utility and min distance of the selected points. GIST noticeably outperforms simple, greedy, and random. The poor performance of greedy is somewhat surprising since it is normally competitive for submodular-style objectives. simple also outperforms greedy for all but the extreme values of $k$, i.e., very small or close to $n$, but is always worse relative to GIST.
5.2 Image classification
The following data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
Setup.
We use the standard vision dataset ImageNet (Russakovsky et al., 2015), which contains ~1.3 million images and 1000 classes. We select 10% of the data points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016). Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example. The uncertainty value of $\mathbf{x}_{i}$ is $u_{i}=1-\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})$, where $b$ is the best class label according to the initial model. Finally, we use the fast maximum inner product search of Guo et al. (2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance. We present all of the model training hyperparameters in Appendix C.
We compare GIST with three state-of-the-art benchmarks:
  • random: We select samples from the dataset uniformly at random (and without replacement). While being a simple and lightweight approach, this implicitly encodes diversity in many settings and provides good solutions.
  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using the margin scores $m_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example.
  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph. The distance between non-adjacent nodes is implicitly assumed to be very large, which is acceptable in the large budget regime $k=\Omega(n)$.
For this task, GIST uses uncertainty values as its weights $w_{i}=u_{i}$ and the same graph as $k$-center, and sets $\alpha=0.85$.
Results.
We run each algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we use to train a new ResNet-56 model. We report the top-1 classification accuracy of these models in Table 1.
$k$ (of data) random margin $k$-center GIST
30 65.02 67.64 67.32 67.70
40 68.92 70.45 70.98 71.20
50 70.86 73.41 72.51 73.67
60 72.39 74.70 74.78 75.16
70 72.92 75.57 75.62 75.85
80 74.36 75.39 76.26 76.49
90 75.10 76.11 76.02 76.50
Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms. The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples.
Conclusion
This work proposes a novel subset selection problem called MDDS that combines the total utility of the selected points with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective. We design and analyze the GIST algorithm, which achieves a $\nicefrac{{2}}{{3}}$-approximation guarantee for general metric spaces by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm. We also give a nearly linear-time implementation of GIST for points in low-dimensional Euclidean space. We complement our algorithm with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$, in general metric spaces. We also prove hardness of approximation in the practical cases of Euclidean metrics or $k=\Omega(n)$. Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDDS (in particular greedy) on a simple set of Gaussian points with random weights. We conclude with an experiment comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet.
Appendix A Missing analysis for Section 3
A.1 Proof of Theorem 3.4
See 3.4
Proof.
First, observe that we can compute and store all pairwise distances in $O(n^{2}d)$ time and $O(n^{2})$ space. This is useful in the naive implementation because we can then look up the values $\textnormal{dist}(u,v)$ in constant time in any subroutine.
In the nearly linear-time implementation, we skip this precomputation step and use cover trees for fast exact nearest-neighbor search in low-dimensional Euclidean space (Beygelzimer et al., 2006): inserting a point into the index takes $O(c^{12}\log n)$ time and finding a nearest neighbor in the index to a query point takes $O(c^{6}\log n)$ time, where $c=\Theta(d)$ is the doubling dimension of $d$-dimensional Euclidean space.
Thus, we can speed up the following subroutines in Algorithm 1 to have near-linear running times in a fixed dimension:
  • Computing $d_{\min}$: $O(n^{2})\rightarrow O(n\cdot d^{12}\log n)$ after discarding duplicate points.
  • GreedyIndependentSet: $O(nk)\rightarrow O(n\cdot d^{12}\log n)$
  • Evaluating $f(S)$: $O(k^{2})\rightarrow O(k\cdot d^{12}\log k)$
Taking advantage of the error $\varepsilon$ in our $(\nicefrac{{2}}{{3}}-\varepsilon)$-approximation guarantee, we use the approximate point-set diameter algorithm in Imanparast et al. (2018, Theorem 1) to compute a $(1-\varepsilon)$-approximate diameter, together with two points realizing it. This speeds up the naive running time of Lines 34 in Algorithm 1 as follows:
  • Computing a $(1-\varepsilon)$-approximate $d_{\max}$: $O(n^{2})\rightarrow O\left(n\cdot 2^{d}d+\frac{(2\sqrt{d})^{2d}}{\varepsilon^{d-1}}\right)$
Since we have a smaller diameter $(1-\varepsilon)d_{\max}$, we need to show that it is compatible with the proof of Theorem 3.1. Concretely, this means modifying Eq. (5) and (6), and observing that the inequality
$\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)(1-\alpha)\cdot(1-\varepsilon)d^{*}\geq\left(\frac{2}{3}-\varepsilon\right)(1-\alpha)d^{*}$
still holds. Lastly, the number of calls to GreedyIndependentSet (i.e., distance thresholds that GIST tries) is $\min\{\lceil\log_{1+\varepsilon}(2/\varepsilon)\rceil,n^{2}\}$. Putting everything together proves the claim. ∎
Appendix B Missing analysis for Section 4
B.1 Proof of Theorem 4.2
See 4.2
Proof.
We use the vertex cover instance constructed in Håstad (2001, Theorem 8.1). For any $\varepsilon_{0}>0$, they construct a graph $G$ with $2^{r+2}$ nodes for some integer $r$ and show that it is NP-hard to distinguish between the following two cases:
  • YES case: Graph $G$ has an independent set of size $2^{r}(1-\varepsilon_{0})$.
  • NO case: Graph $G$ has no independent set of size larger than $2^{r}(\nicefrac{{1}}{{2}}+\varepsilon_{0})$.
We mimic the general structure of the proof of Theorem 4.1. Add a point for each vertex in graph $G$ and set its weight to $1$. Set the distance between two points to $1$ if they share an edge in the graph, and set it to $2$ otherwise. Lastly, let $k=2^{r}$. Since there are $2^{r+2}$ nodes in the graph, $k=|V|/4$ in this case. In the YES case, objective function (1) can achieve the following value by selecting the independent set of size $2^{r}(1-\varepsilon_{0})$:
OPT $\geq\alpha\cdot\frac{1}{2^{r}}\cdot(2^{r}(1-\varepsilon_{0}))+(1-\alpha)\cdot 2$
$=2-(1+\varepsilon_{0})\cdot\alpha.$
In the NO case, the largest value that can be achieved is the best of two scenarios: (a) selecting $k$ points with minimum distance $1$, or (b) selecting an independent set of size at most $2^{r}(\nicefrac{{1}}{{2}}+\varepsilon_{0})$ and achieving minimum distance $2$. This results in the following upper bound for the objective:
ALG $\leq\max\left\{\alpha+(1-\alpha)\cdot 1,\alpha\cdot(\nicefrac{{1}}{{2}}+\varepsilon_{0})+(1-\alpha)\cdot 2\right\}$
$=\max\left\{1,2-(\nicefrac{{3}}{{2}}-\varepsilon_{0})\cdot\alpha\right\}.$
The largest hardness gap occurs when the two terms of the $\max$ expression become equal, which happens at $\alpha=1/(\nicefrac{{3}}{{2}}-\varepsilon_{0})$ and yields the upper bound $\textnormal{ALG}\leq 1$. The optimum value OPT will be lower bounded by:
OPT $\geq 2-(1+\varepsilon_{0})\cdot\alpha$
$=2-\frac{1+\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}$
$=\frac{2-3\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}.$
The desired inapproximability gap of $\nicefrac{{3}}{{4}}+\varepsilon$ in Theorem 4.2 is achieved when the ratio between the upper bound on ALG and lower bound for OPT drops below it. In other words, we want to set $\varepsilon_{0}$ such that the following inequality holds:
$\frac{2-3\varepsilon_{0}}{\nicefrac{{3}}{{2}}-\varepsilon_{0}}>\frac{1}{\nicefrac{{3}}{{4}}+\varepsilon}.$
We set $\varepsilon_{0}\coloneqq\varepsilon$ and prove that the above inequality holds. We first note that both the numerator and denominator of the left side ratio are positive. Therefore, we can invert both ratios and swap their places to achieve the following equivalent inequality:
$\frac{3}{4}+\varepsilon>\frac{\nicefrac{{3}}{{2}}-\varepsilon_{0}}{2-3\varepsilon_{0}}\iff\varepsilon>\frac{\nicefrac{{3}}{{2}}-\varepsilon_{0}}{2-3\varepsilon_{0}}-\frac{3}{4}=\frac{6-4\varepsilon_{0}-6+9\varepsilon_{0}}{8-12\varepsilon_{0}}>\frac{5\varepsilon_{0}}{8}.$
The left-hand side inequality $\varepsilon>\nicefrac{{5\varepsilon_{0}}}{{8}}$ holds by our choice of $\varepsilon_{0}=\varepsilon$. ∎
B.2 Proof of Lemma 4.4
@@ Appendix B Missing analysis for Section 4
First, let $n=|V|$ and $m=|E|$. Augment $G$ by adding a self-loop to each node to get $G^{\prime}=(V,E^{\prime})$. We embed $V$ using $G^{\prime}$ since each node now has positive degree. Let $\deg^{\prime}(v)$ be the degree of $v$ in $G^{\prime}$ and $N^{\prime}(v)$ be the neighborhood of $v$ in $G^{\prime}$.
@@
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$
$=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
@@
We reduce this instance of bounded-degree maximum independent set to MDDS with objective function (1) as follows. Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4. We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=1$. In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (1):
@@
Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a positive constant. This completes the proof of APX-completeness. ∎
Appendix C Missing details for Section 5

### v2

@@ Appendix A $f(S)$ is Not Submodular
See 3.4
@@
Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$. There are two cases: $d^{*}<\varepsilon d_{\max}$ and $d^{*}\geq\varepsilon d_{\max}$. In the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation. To see this, first observe that
$\textnormal{OPT}\geq d_{\max}>\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>d^{*}.$
The sum of the $k$ heaviest points upper bounds $g(S^{*})$, so we have
$\textnormal{ALG}\geq g(S^{*})=\textnormal{OPT}-d^{*}>(1-\varepsilon)\textnormal{OPT}.$
Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
@@
$f(T)$ $\geq g(S^{*})+d\geq g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}.$ (5)
@@
$\textnormal{ALG}\geq d_{\max}\geq d^{*}.$ (6)
Putting (5) and (6) together, for any $0\leq p\leq 1$, it holds that
ALG $\geq p\cdot\left(g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}\right)+(1-p)\cdot d^{*}$
$=p\cdot g(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot d^{*}.$
To maximize the approximation ratio as $\varepsilon\rightarrow 0$, we set $p=2/3$ by solving $p=1-p/2$. It follows that
ALG $\geq\frac{2}{3}\cdot g(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot d^{*}$ (7)
$=\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot d^{*}$
$\geq\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot d^{*}$
$\geq\left(\nicefrac{{2}}{{3}}-\varepsilon\right)\cdot\textnormal{OPT},$
@@
Appendix C Missing Analysis for Section 4
C.1 Proof of Theorem: 4.2
See 4.2
@@ Appendix C Missing Analysis for Section 4
We reduce this instance of the maximum clique decision problem to MDMS with objective function (2) as follows. Represent each vertex of graph $G$ with a point in our ground set. The distance between a pair of points is $2$ if there is an edge between their corresponding vertices in $G$, and it is $1$ otherwise.
@@
C.2 Proof of Lemma 4.4
@@
Let $n=|V|$ and $m=|E|$. Augment $G$ by adding a self-loop to each node to get $G^{\prime}=(V,E^{\prime})$. We embed $V$ using $G^{\prime}$ since each node now has positive degree. Let $\deg^{\prime}(v)$ be the degree of $v$ in $G^{\prime}$ and $N^{\prime}(v)$ be the neighborhood of $v$ in $G^{\prime}$.
@@
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
@@
We reduce this instance of bounded-degree maximum independent set to MDMS with objective function (2) as follows. Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4. We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=1$. In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (2):
@@
Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a constant. This completes the proof of APX-completeness. ∎
Appendix D Additional Details for Section 5
D.1 Synthetic Dataset
We extend our comparison of baseline algorithms in Section 5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
Figure 2: Baseline comparison with $\alpha\in(0.85,0.90,0.95,1.00)$ and $\beta=0.75$.
Figure 3: Baseline comparison with $\alpha=0.95$ and $\beta\in(0.60,0.70,0.80,0.90)$.
D.2 Image Classification